A man will be three times as old as his son will be in nine years. If he is six times as old as his son is now, how old is the SON? .
What I have come up with is as follows.
3n+9 = 6n
But I am not positvie this equations even works, thanks for the help.

Answer by scott8148(5895)

let m="man now" and s="son now"

"A man will be three times as old as his son will be in nine years" ___ m+9=3(s+9)

"he is six times as old as his son is now" ___ m=6s

"how old is the SON?" ___ substituting ___ (6s)+9=3(s+9)

expanding ___ 6s+9=3s+27

subtracting 3s+9 ___ 3s=18

dividing by 3 ___ s=6